# BatchLLM: Optimizing Large Batched LLM Inference with Global Prefix Sharing and Throughput-oriented Token Batching

### v1

@@ Abstract.
Large language models (LLMs) increasingly play an important role in a wide range of information processing and management tasks. Many of these tasks are performed in large batches or even offline, and the performance indictor for which is throughput. These tasks usually show the characteristic of prefix sharing, where different prompt input can partially show the common prefix. However, the existing LLM inference engines tend to optimize the streaming requests and show limitations of supporting the large batched tasks with the prefix sharing characteristic. The existing solutions use the LRU-based cache to reuse the KV context of common prefix between requests. The KV context that are about to be reused may prematurely evicted with the implicit cache management. Even if not evicted, the lifetime of the shared KV context is extended since requests sharing the same context are not scheduled together, resulting in larger memory usage. These streaming oriented systems schedule the requests in the first-come-first-serve or similar order. As a result, the requests with larger ratio of decoding steps may be scheduled too late to be able to mix with the prefill chunks to increase the hardware utilization. Besides, the token and request number based batching can limit the size of token-batch, which keeps the GPU from saturating for the iterations dominated by decoding tokens. We propose BatchLLM to address the above problems. BatchLLM explicitly identifies the common prefixes globally. The requests sharing the same prefix will be scheduled together to reuse the KV context the best, which also shrinks the lifetime of common KV memory. BatchLLM reorders the requests and schedules the requests with larger ratio of decoding first to better mix the decoding tokens with the latter prefill chunks, and applies memory-centric token batching to enlarge the token-batch sizes, which helps to increase the GPU utilization. Finally, BatchLLM optimizes the prefix-shared Attention kernel with horizontal fusion to reduce tail effect and kernel launch overhead. Extensive evaluation shows that BatchLLM outperforms vLLM by $\times1$ to 2.0×2.0$\times$2.0 × on a set of microbenchmarks and two typical industry workloads.
@@ 1. Introduction
Given the limitations of existing works, the LLM inference engines face significant demands in supporting emerging throughput-oriented large batched tasks, especially as the LLM is more and more widely used in the huge amounts of batched information processing and management.
@@
Building upon these insights, this paper proposes a set of optimizations. It builds the global prefix tree on the large batch ahead-of-time to enable the explicit prefix sharing and avoid the problem of premature eviction of KV contexts. To simplify the design and reduce the overhead of token batching and Attention kernel on multi-level prefix, it simplifies the multi-level prefix into a single level one while retaining the similar KV reusing ratio with Dynamic Programming algorithm on the tree. It schedules the requests sharing the common prefix together to enable them reusing the KV context and reduce the lifetime of the common KV. To form large token-batches at each iteration to increase GPU utilization, it reorders the requests ahead-of-time according to the length of prompt and the decoding (estimated). The request with larger ratio of decoding length to prompt length will be scheduled first so that the decoding tokens will be processed earlier and can be mixed with the latter longer prefill chunks. It also batches the tokens according to the KV memory usage rather than only depending on request number threshold, to form large batches for decoding tokens. Finally, it applies horizontal fusion of Attention calculation on different KV chunks to reduce the overhead of tail effect and kernel launch of the basic prefix-shared Attention implementation.
We have implemented BatchLLM on top of vLLM (Kwon et al., 2023) and conducted extensive evaluations using a set of microbenchmarks of different datasets and models and two industry workloads on NVIDIA and AMD GPUs. BatchLLM achieves end-to-end performance improvement of $\times1$ to 2.0×2.0$\times$2.0 × than vLLM for the microbenchmarks and the two industry workloads. The performance gain comes from BatchLLM’s ability to reuse the common prefix explicit in the global view, form the token-batches with higher compute-intensity (the ratio of computational operations to memory size accessed), and the more efficient prefix-shared Attention.
@@ 2. LLM Inference Background
2.1. LLM Inference and Its Performance Factors
The auto-regressive LLM inference usually consists of two phases: the prefill phase to process the input prompt, and the decoding phase to generate the output.
The prefill phase is usually compute-intensive and bounded by the massive computation. This is because the input tokens are all known ahead, and the computation of these tokens can be performed in parallel. The decoding phase is performed differently from the prefill phase: the output tokens are generated one by one and cannot be performed in parallel333 In this paper, we mainly focus on the basic decoding scheme and will not discuss the parallel decoding technique (Cai et al., 2024).. As a result, the decoding phase can be easily bounded by the memory access of the weight due to the memory wall problem (Xia et al., 2023, 2024). Enlarging the batch size can help to push the decoding computation to compute-bound, but the batch size may be limited by the GPU memory capacity. In real practice, the batching of the decoding can be suboptimal due to either the lack of memory space or the inefficient token batching, leading to low hardware utilization.
Attention can be an important performance factor when the sequence is long. According to its algorithm pattern, the basic Attention computation in decoding phase is the matrix-vector multiplication between the (batched) vector $Q$ and matrix $KV$ containing all history context, as tokens from different requests have different KV context. Due to this characteristic, enlarging the batch size usually can not increase the compute utilization of the GPU as it does not increase the compute-intensity.
@@ 2.2. KV Cache and Prefix Sharing
Based on the nature of Attention calculation (Vaswani et al., 2017), the processing of each token (in both prefill and decoding phase) will generate the corresponding KV context. The computation of latter tokens consume the KV context of earlier tokens to perform Attention calculation. During the processing of a request (i.e., a prompt), the KV context is usually cached in memory for each of the processed token (called KV cache) so that the latter tokens can reuse the history KV context in memory for Attention calculation, rather than recomputing the history KV.
The KV cache can consume a large amount of GPU memory, which can limit the batch size of LLM inference and thus hurt the GPU hardware utilization of the linear layers consist of matrix multiplication (MatMul) calculation. The KV memory size of a single token is $2\times n_{bytes}\times n_{layers}\times n_{kvhead}\times d_{head}$ Bytes. Take Mistral 7B model (Jiang et al., 2023) with grouped-query attention (Ainslie et al., 2023) as an example, which is a small size LLM, each token takes 0.125MiB memory if it uses float16 datatype for KV, and a sequence of 2K length will take 256MiB memory for the KV cache. An NVIDIA A10 GPU with 24GiB memory can only hold less than 100 such sequences’ KV memory, even without considering the memory consumption of the weight itself. The KV memory is usually managed in blocked memory to reduce memory fragment and enable KV sharing (Kwon et al., 2023), the logical continuous KV tensors are mapped to the physical discrete memory blocks mapped with the block table.
@@ 2.3. Per-iteration Token Batching
Due to the dynamic output length of LLM execution, batching the tasks at the granularity of request results in a waste of the GPU resource, as some requests may finish earlier. Instead, the current LLM inference engines batche the tasks at the granularity of tokens. Specifically, it forms the token-batch at each iteration and allows to add new tokens into the batch dynamically (Yu et al., 2022), which is also called continuous batching. During the execution, the linear operators of different tokens will be concatenated together into a larger linear operator for execution at each iteration.
@@ 3. Opportunities and Insight
3.1. Emerging Scenarios: Large Batched Through Oriented Tasks with Shared Prefix
LLM has been increasingly used in a broad range of information processing and management tasks (Wang et al., 2023; Xiong et al., 2024; Meguellati et al., 2024; Zhao et al., 2024; Kim et al., 2024; Chen et al., 2023; Fernandez et al., 2023). Due to the large traffic of these tasks and the high cost of LLM inference, many of these tasks are processed in very large batch (or even offline) whose performance metric is mainly the throughput rather than latency. Take the snippet generation task (Wang et al., 2023; Xiong et al., 2024) in search engine as an example, which is to generate the snippet of the web document (web content) in the search result page according to both document and user query. An example prompt in a previous work (Wang et al., 2023) is: generate a short snippet based on the given Document to answer the given Query. Document: <…> Query: <…>. Due to that it is hard to meet the SLO when performing the massive LLM inferences online, a common practice is to run the snippet generation task offline for the high-frequency web documents and queries, and retrieve the offline results during online serving when the corresponding document-query pair occurs.
Many of the LLM-based information processing and management tasks show the characteristic of the shared prefix between the prompts of different requests, which comes from the nature that different tasks can perform on the same content (e.g., web document). Take the search engine task as an example, it may extract different information from the same web document so that the document can be a shared prefix in the prompt. Specifically, as for the snippet generation task, given that it is to extract the snippet of the document-query pair, the document can be the common prefix if it appears for different queries.
The shared prefix can be managed in multiple levels. For example, the first level prefix can be the global system information and instructions, and the second level can be the web document. However, with the development of LLM and the using of task-specific model fine-tuning, the prefix of system information and instructions becomes shorter and shorter, and the length of the prefix will finally be dominated by the web document itself. On the one hand, the LLM models are incorporating more and more reasoning abilities into themselves (like OpenAI o1 (OpenAI, 2024a)) and will not rely on complex prompts. On the other hand, the task-specific fine-tuning can help to incorporate the complex prompts into the model weight. As a result, the prefix sharing of the long context rather than the other instructions can be the most important for many of the information processing and management tasks. This motivates the first-level prefix enlargement in Sec.4.2.
@@ 3.2. The New Optimization Demands
1. The implicit prefix caching cannot lead to the optimal KV reuse for the large batched inference. The existing LLM serving systems (Kwon et al., 2023; Zheng et al., 2023) use the prefix tree to maintain the KV blocks. The same prefix (identified by runtime hashing) can be mapped to the same KV blocks to avoid repeated computation. It uses the LRU policy to evict the blocks from the block table. As for a large batch of inputs to be processed, it does not keep the prefix in memory in the global view in the large batch, but only according to the history input. As a result, the shareable KV blocks can be easily evicted prematurely. We have evaluated vLLM’s prefix sharing with a typical industry workload. The optimal token saving ratio by prefix reusing is 56% by analyzing the input dataset. Note that the saving ratio is calculated by:
@@
$N_{processed\_prefill\_tokens}$ is the number of prefill tokens processed after prefix sharing, and $N_{logical\_prefill\_tokens}$ is the original prefill token number of all the requests. The vLLM’s implicit prefix caching only achieves 44.2% (refer to Sec.6.3.1).
There is a demand to enable better prefix sharing for the large batched tasks, especially considering that all the prefill characteristics in the batch are known ahead.
@@
2. The online serving oriented scheduling and token batching can be suboptimal for the prefix-shared and throughput oriented tasks. The current LLM inference systems schedule the tasks at the granularity of request. They do not analyze the prefix sharing characteristics of the whole large batch nor schedule the requests sharing common prefix together. This not only can evict the shareable KV context prematurely as discussed above, but also extends the lifetime of the shareable KV memory, exacerbating the problem of large KV memory consumption. Besides, these systems design tends to support online services. They form the token-batches at each iteration with the constraint of requests’ arriving order, which can result in suboptimal token-batches. Take the example in Fig.1, the naive scheduling in the order of request’s coming cannot mix the decoding of request 3 with the prefill chunks of other requests (chunked-prefill background in Sec.2.3). Instead, by scheduling request 3 earlier, its decoding can be mixed with the prefill chunks of other requests. Another problem is that, the current systems use the token and request number in the token-batch as the threshold to limit the batching, which limits to batch more tokens together in the decoding dominated token-batches and prevents better utilizing the GPU even when the memory is enough. Fig.2 shows the number of tokens at each iteration for a typical industry task (Sec.6.2.3) performing with vLLM’s best configuration. It shows that there are "valleys"444 This paper uses valley to express the token-batches that has smaller number of tokens, appearing as the valley in the timeline of token numbers. at many iterations where the number of tokens may not be large enough to saturate the GPU.
There is a demand to reschedule the tasks to make the requests sharing common prefix closer, and enable better mix of decoding tokens with prefill chunks. There is also a room to enlarge the size of token-batches dominated by the decoding tokens.
@@
3. There is still room for Attention optimization of prefix sharing. As described in Sec.2.2, the existing works (Ye et al., 2024b; Zhu et al., 2024; Juravsky et al., 2024; Ye et al., 2024a) support prefix-shared Attention by computing the Attention on different chunks of KV, performing MatMul on the common prefix and the basic matrix-vector multiplication on the non-shared chunk. But the computation on different chunks are in different kernels. On the one hand, the separated kernels increase the tail effect of GPU kernel. On the other hand, the multiple kernels
@@ 3.3. Insight
To optimize the throughput oriented tasks described in Sec.3.1 and address the limitations described in Sec.3.2, by leveraging the characteristics of the large batch, BatchLLM has the following insights.
1. Explicit prefix identification and sharing. Instead of hashing and matching the common prefix at runtime and manage the prefix with LRU cache, BatchLLM explicitly identifies the common prefix globally within the large batch ahead-of-time, which will not miss the opportunity of prefix sharing caused by the implicit cache. Besides, BatchLLM refactors the prefix tree with a Dynamic Programming algorithm to enlarge the first-level prefix to avoid the system complexity and kernel overhead of the multi-level prefix. Details are described in Sec.4.2.
@@
3. Attention kernel optimization with horizontal fusion. BatchLLM horizontally fuses the computation on different KV chunks into the same kernel to reduce the tail effect and kernel launch overhead. This method is not novel, but effective.
@@ 4.1. Overview
Fig.3 shows the overview of BatchLLM optimizations. First, it analyzes the large batch of input prompts and identifies the common prefix explicitly. This is done before the scheduling of the requests. It simplifies the multi-level prefix into a single level prefix with the insight that the prefixes of the information processing and management are usually dominated by a long context (e.g., web document), as discussed in Sec.3.1. The explicit prefix processing is illustrated in Sec.4.2. The requests will be organized into groups where each group corresponds to the quests sharing the same prefix. The group will be the basic unit of task scheduling. Before scheduling the groups, BatchLLM will also reorder the groups according to the ratio of prefill length and estimated decoding length and postpone the groups with longer prefill and shorter decoding. It then forms the token-batches with the consideration of the KV memory usage. This aims to better mix the decoding steps with the prefill chunks to increase the overall token-batch size. The throughput-oriented scheduling and token-batching optimization is described in Sec.4.3. BatchLLM also implements the horizontal fused Attention kernel to reduce the tail effect and kernel launch overhead, described in Sec.4.4. We have integrated the above optimization into the state-of-the-art LLM inference engine (vLLM (Kwon et al., 2023)).

### v2

@@ Abstract.
Large language models (LLMs) increasingly play an important role in a wide range of information processing and management tasks. Many of these tasks are performed in large batches or even offline, and the performance indictor for which is throughput. These tasks usually show the characteristic of prefix sharing, where different prompt input can partially show the common prefix. However, the existing LLM inference engines tend to optimize the streaming requests and show limitations of supporting the large batched tasks with the prefix sharing characteristic. The existing solutions use the LRU-based cache to reuse the KV context of common prefix between requests. The KV context that are about to be reused may prematurely evicted with the implicit cache management. Besides, the streaming oriented systems do not leverage the request-batch information and can not mix the decoding tokens with the prefill chunks to the best for the batched scenarios, and thus fails to saturate the GPU. We propose BatchLLM to address the above problems. BatchLLM explicitly identifies the common prefixes globally. The requests sharing the same prefix will be scheduled together to reuse the KV context the best. BatchLLM reorders the requests and schedules the requests with larger ratio of decoding first to better mix the decoding tokens with the latter prefill chunks, and applies memory-centric token batching to enlarge the token-batch sizes, which helps to increase the GPU utilization. Finally, BatchLLM optimizes the prefix-shared Attention kernel with horizontal fusion to reduce tail effect and kernel launch overhead. Extensive evaluation shows that BatchLLM outperforms vLLM and SGLang by 1.3×1.3$\times$1.3 × to 10.8×10.8$\times$10.8 × on a set of microbenchmarks and a typical industry workload under different hardware environments.
@@ 1. Introduction
Building upon these insights, this paper proposes a set of optimizations. It builds the global prefix tree on the large batch ahead-of-time to enable the explicit prefix sharing and avoid the problem of premature eviction of KV contexts. To simplify the design and reduce the overhead of token batching and Attention kernel on multi-level prefix, it simplifies the multi-level prefix into a single level one while retaining the similar KV reusing ratio with Dynamic Programming algorithm on the tree. It schedules the requests sharing the common prefix together to enable them reusing the KV context. To form large token-batches at each iteration to increase GPU utilization, it reorders the requests ahead-of-time according to the length of prompt and the decoding (estimated). The request with larger ratio of decoding length to prompt length will be scheduled first so that the decoding tokens will be processed earlier and can be mixed with the latter longer prefill chunks. It also batches the tokens according to the KV memory usage rather than only depending on request number threshold, to form large batches for decoding tokens. Finally, it applies horizontal fusion of Attention calculation on different KV chunks to reduce the overhead of tail effect and kernel launch of the basic prefix-shared Attention implementation.
We have implemented BatchLLM on top of vLLM (Kwon et al., 2023) and conducted extensive evaluations using a set of microbenchmarks of different datasets and models and an industry workload on NVIDIA and AMD GPUs. BatchLLM achieves end-to-end performance improvement of 1.3×1.3$\times$1.3 × to 10.8×10.8$\times$10.8 × than vLLM and SGLang for the microbenchmarks and an industry workload under different hardware environments. The performance gain comes from BatchLLM’s ability to reuse the common prefix explicit in the global view, form the token-batches with higher compute-intensity, and the more efficient prefix-shared Attention.
@@ 2. LLM Inference Background
2.1. Performance Factors of LLM Inference
The auto-regressive LLM inference usually consists of two phases: the prefill phase to process the input prompt, and the decoding phase to generate the output. The prefill phase is usually compute-intensive and bounded by the massive computation. This is because the input tokens are all known ahead, and the computation of these tokens can be performed in parallel. In the decoding phase, the output tokens are generated one by one and cannot be performed in parallel333 In this paper, we mainly focus on the basic decoding scheme and will not discuss the parallel decoding technique (Cai et al., 2024).. As a result, the decoding phase can be easily bounded by the memory access of the weight due to the memory wall problem (Xia et al., 2023, 2024). Enlarging the batch size can help to push the decoding computation to compute-bound, but the batch size may be limited by the GPU memory capacity. In real practice, the batching of the decoding can be suboptimal due to either the lack of memory space or the inefficient token batching, leading to low hardware utilization. Attention can be an important performance factor when the sequence is long. According to its algorithm pattern, the basic Attention computation in decoding phase is the matrix-vector multiplication between the (batched) vector $Q$ and matrix $KV$ containing all history context, as tokens from different requests have different KV context. Due to this characteristic, enlarging the batch size usually can not increase the compute utilization of the GPU as it does not increase the compute-intensity.
@@ 2.2. KV Cache and Prefix Sharing
Based on the nature of Attention calculation (Vaswani et al., 2017), the processing of each token (in both prefill and decoding phase) will generate the corresponding KV context. The computation of latter tokens consume the KV context of earlier tokens to perform Attention calculation. During the processing of a request (i.e., a prompt), the KV context is usually cached in memory for each of the processed token (called KV cache) so that the latter tokens can reuse the history KV context in memory for Attention calculation. The KV cache can consume a large amount of GPU memory, which can limit the batch size of LLM inference and thus hurt the GPU hardware utilization of the linear layers consist of matrix multiplication (MatMul) calculation. The KV cache is usually managed in blocked memory to reduce fragmentation and enable KV sharing (Kwon et al., 2023), the logical continuous KV tensors are mapped to the physical discrete memory blocks with the block table.
@@ 2.3. Per-iteration Token Batching
Due to the dynamic output length of LLM execution, batching the tasks at the granularity of request results in a waste of the GPU resource, as some requests may finish earlier. Instead, the current LLM inference engines batch the tasks at the granularity of tokens. Specifically, it forms the token-batch at each iteration and allows to add new tokens into the batch dynamically (Yu et al., 2022), which is also called continuous batching. During the execution, the linear operators of different tokens will be concatenated together into a larger linear operator for execution at each iteration.
@@ 3. Opportunities and Insight
3.1. Emerging Industry Scenarios
LLM has been increasingly used in a broad range of industry tasks with its in-context learning and reasoning ability (Wang et al., 2023; Xiong et al., 2024; Meguellati et al., 2024; Zhao et al., 2024; Kim et al., 2024; Chen et al., 2023; Fernandez et al., 2023). Due to the large traffic of these tasks and the high cost of LLM inference, many of these tasks are processed in very large batch (or even offline) whose performance metric is mainly the throughput rather than latency. Take the snippet generation task (Wang et al., 2023; Xiong et al., 2024) in search engine as an example, which is to generate the snippet of the web document (web content) in the search result page according to both document and user query. An example prompt in a previous work (Wang et al., 2023) is: generate a short snippet based on the given Document to answer the given Query. Document: <…> Query: <…>. Due to that it is hard to meet the SLO when performing the massive LLM inferences online, a common practice is to run the snippet generation task offline for the high-frequency web documents and queries, and retrieve the offline results during online serving when the corresponding document-query pair occurs.
Many of the LLM-based information processing and management tasks show the characteristic of the shared prefix between the prompts of different requests, which comes from the nature that different tasks can perform on the same content (e.g., web document). Take the search engine task as an example, it may extract different information from the same web document so that the document can be a shared prefix in the prompt. Specifically, as for the snippet generation task, given that it is to extract the snippet of the document-query pair, a document can be the common prefix for different queries.
The shared prefix can be managed in multiple levels. For example, the first level prefix can be the global system information and instructions, and the second level can be the web document. However, with the development of LLM and the using of task-specific model fine-tuning, the prefix of system information and instructions becomes shorter and shorter, and the length of the prefix will finally be dominated by the web document itself. On the one hand, the LLM models are incorporating more and more reasoning abilities into themselves (like OpenAI o1 (OpenAI, 2024a)) and will not rely on complex prompts. On the other hand, the task-specific fine-tuning can help to incorporate the complex prompts into the model weight. As a result, the prefix sharing of the long context rather than the instructions can be the most important for many industry tasks. This motivates the first-level prefix enlargement in Sec.4.2.
@@ 3.2. The New Optimization Demands
1. The implicit prefix caching cannot lead to the optimal KV reuse for the large batched inference. The existing LLM serving systems (Kwon et al., 2023; Zheng et al., 2023) use the prefix tree to maintain the KV blocks. The same prefix (identified by runtime hashing) can be mapped to the same KV blocks to avoid repeated computation. It uses the LRU policy to evict the blocks from the block table. As for a large batch of inputs to be processed, it does not keep the prefix in memory in the global view in the large batch, but only according to the history input. As a result, the shareable KV blocks can be easily evicted prematurely. We have evaluated vLLM’s prefix sharing with a typical industry workload. The optimal token saving ratio by prefix reusing is 58.1% by analyzing the input dataset. Note that the saving ratio is calculated by:
@@
$N_{processed\_prefill\_tokens}$ is the number of prefill tokens processed after prefix sharing, and $N_{logical\_prefill\_tokens}$ is the original prefill token number of all the requests. The vLLM’s implicit prefix caching only achieves 35.8% (refer to Sec.6.3.1). There is a demand to enable better prefix sharing for the large batched tasks, especially considering that all the prefill characteristics in the batch are known ahead.
@@
2. The online serving oriented scheduling and token batching can be suboptimal for the prefix-shared and throughput oriented tasks. The current LLM inference systems schedule the tasks at the granularity of request. They do not analyze the prefix sharing characteristics of the whole large batch nor schedule the requests sharing common prefix together. This not only can evict the shareable KV context prematurely as discussed above, but also extends the lifetime of the shareable KV memory, exacerbating the problem of large KV memory consumption. Besides, these systems design tends to support online services. They form the token-batches at each iteration with the constraint of requests’ arriving order, which can result in suboptimal token-batches. Take the example in Fig.1, the naive scheduling in the order of request’s coming cannot mix the decoding of request 3 with the prefill chunks of other requests (chunked-prefill background in Sec.2.3). Instead, by scheduling request 3 earlier, its decoding can be mixed with the prefill chunks of other requests. Another problem is that, the current systems use the token and request number in the token-batch as the threshold to limit the batching, which limits to batch more tokens together in the decoding dominated token-batches and prevents better utilizing the GPU even when the memory is enough. Fig.2 shows the number of tokens at each iteration for a typical industry task (Sec.6.2.2) performing with vLLM’s best configuration. It shows that there are "valleys"444 This paper uses valley to express the token-batches that has smaller number of tokens, appearing as the valley in the timeline of token numbers. at many iterations where the number of tokens are not large enough to saturate the GPU. There is a demand to reschedule the tasks to make the requests sharing common prefix closer, and enable better mix of decoding tokens with prefill chunks. There is also a room to enlarge the size of token-batches dominated by the decoding tokens.
@@
3. There is still room for Attention optimization of prefix sharing. As described in Sec.2.2, the existing works (Ye et al., 2024b; Zhu et al., 2024; Juravsky et al., 2024; Ye et al., 2024a) support prefix-shared Attention by computing the Attention on different chunks of KV, performing MatMul on the common prefix and the basic matrix-vector multiplication on the non-shared chunk. But the computation on different chunks are in different kernels. On the one hand, the separated kernels increase the tail effect of GPU kernel. On the other hand, the multiple kernels introduce extra launching overhead.
@@ 3.3. Insight
BatchLLM has the following insights to address the limitations described in Sec.3.2.
1. Explicit prefix identification and sharing. Instead of hashing and matching the common prefix at runtime and manage the prefix with LRU cache, BatchLLM explicitly identifies the common prefix globally within the large batch ahead-of-time, which will not miss the opportunity of prefix sharing caused by the implicit cache. Besides, BatchLLM refactors the prefix tree with a Dynamic Programming algorithm to enlarge the first-level prefix to avoid the system complexity and kernel overhead of the multi-level prefix.
@@
3. Attention kernel optimization with horizontal fusion. BatchLLM horizontally fuses the computation on different KV chunks into the same kernel to reduce the tail effect and kernel launch overhead.
@@ 4.1. Overview
Fig.3 shows the overview of BatchLLM optimizations. First, it analyzes the large batch of input prompts and identifies the common prefix explicitly. This is done before the scheduling of the requests. It simplifies the multi-level prefix into a single level prefix with the insight that the prefixes of the industry tasks are usually dominated by a long context (e.g., web document), as discussed in Sec.3.1. The explicit prefix processing is illustrated in Sec.4.2. The requests will be organized into groups where each group corresponds to the queries sharing the same prefix. The group will be the basic unit of task scheduling. Before scheduling the groups, BatchLLM will also reorder the groups according to the ratio of prefill length and postpone the groups with longer prefill. It then forms the token-batches with the consideration of the KV memory usage. This aims to better mix the decoding steps with the prefill chunks to increase the overall token-batch size. The throughput-oriented scheduling and token-batching optimization is described in Sec.4.3. BatchLLM also implements the horizontal fused Attention kernel to reduce the tail effect and kernel launch overhead, described in Sec.5.